OpenStudy (pedro):

Lyle has a fourteen-cup box of cereal. If one serving is 3/4 cup, how many servings are in the box?

OpenStudy (anonymous):

\[14/.75=18\frac{2}{3} \]

OpenStudy (anonymous):

\[14\div\frac{3}{4}=14\times \frac{4}{3}=\frac{56}{3}=18\tfrac{2}{3}\]